OpenStudy (anonymous):

whats the translation?? \[g(x)=\sqrt{x-4}\]

OpenStudy (tkhunny):

As opposed to \(t(x) = \sqrt{x}\)?

OpenStudy (anonymous):

\[f(x)=\sqrt{x}\]

OpenStudy (anonymous):

would it be a vertical translation 4 units up?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

compared to \(f(x)\), the graph of \(f(x-4)\) is shifted to the RIGHT 4 units

OpenStudy (anonymous):

if you want to shift \(y=\sqrt{x}\) up by 4 units you would put \(y=\sqrt{x}+4\)

OpenStudy (anonymous):

so a horizontal translation 4 units right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what if the s(x) was now changed to g(x)?